Please help me with these two problems:
(9^-3)/(3^-8)
and
one divided by the cubed root of x to the fourth power?

Answer by littlewiseowl_1997(2)

Hi!
In regards to your first question: working with negative exponents is easier than it looks. Positive exponents mean how many times you multiply by a number; negative exponents mean how many times you DIVIDE by a number. The easiest way to solve these is to write 1 over what the answer would be if the exponent was positive:
(9^-3)= 1/9^3= 1/729 = 0.001371742
(3^-8)= 1/3^8 = 1/6561 = 0.000152416
Now to divide the two: (1/729)/(1/6561)= 9
Now, for the second question:
A cubed root is the same thing as a (1/3) exponent. You can rewrite the problem as such to make it easier.
I need more clarification about how this problem is written: based on how you worded it (without parentheses as a guide), it can be interpreted in a few different ways--- if it is (1/the cubed root of x)^4, or 1/(cubed root of x)^4), for example. If this is clarified, I will try my best in helping you come up with an answer.
Hope this helped a bit!
--- Shannon